01 December 1997
The next ERS-1 check out will take place from 16 to 18 December 1997 covering orbit range 33576 - 33617.
27 November 1997
Synergy of the ATSR-2 and GOME instruments on the detection of biomass burning over Indonesia - ATSR-2 night measurements are compared to GOME nitrogen dioxide measurements over Borneo during autumn 1997.
17 November 1997
Due to the degradation on the GOME PtCrNe-lamp, some problems with the wavelength calibration for the GOME channel 3 have been experienced in level 2 products between 25/10/97 through 14/10/97. The situation has been recovered, the products concerned regenerated and replaced on the ftp server.
15 November 1997
See new images derived from the SAR instrument on board the ERS-2.
07 November 1997
See new night time images derived from the ATSR nadir view calibrated data.
29 October 1997
Please note that besides being constantly updated on this server in the Altimeter Performance section, the USO/SPTR correction files for the ERS-1 and ERS-2 Radar Altimeters are now available at the following addressý
22 October 1997
See the most recent El Nino effect on water temperatures and on drought in Indonesia (images of fires) from the ATSR.
17 October 1997
Full Proceedings of the 3rd ERS 1997 Symposium now available.
07 October 1997
Two maps showing the 23.8 and 36.5 GHz brightness temperatures measured over the South Pole have been added to the Microwave Sounder Performance monitoring service.
Showing 4,671 - 4,680 of 4,884 results.
EO Weekly Newsletter